Question : Computer Network

Question 1 (Short Answer/Essay)

Suppose we want to divide a class-C, 200.0.0.0/24 network into five subnets of the following nominal sizes:

  • A: 64=26 hosts

  • B: 32=25 hosts

  • C: 32=25 hosts

  • D: 32=25 hosts

  • E: 16=24 hosts

Propose subnet-bit values to be assigned for each network. The subnets should not “overlap”.

Copy the table below to your answer box, using richtext mode.

subnet

# host bits needed

# net bits

in last byte

total net bits

subnet bits (each entry 0 or 1)

A

6

8-6=2

/26

___ ___

B

5

8-5=3

/27

___ ___ ___

C

5

8-5=3

/27

___ ___ ___

D

5

8-5=3

/27

___ ___ ___

E

4

8-4=4

/28

___ ___ ___ ___

please provide correct answer
Answer #1

subnet

# host bits needed

# net bits

in last byte

total net bits

subnet bits (each entry 0 or 1)

A

6

8-6=2

/26

11111111.11111111.11111111.11000000

B

5

8-5=3

/27

11111111.11111111.11111111.11100000

C

5

8-5=3

/27

11111111.11111111.11111111.11100000

D

5

8-5=3

/27

11111111.11111111.11111111.11100000

E

4

8-4=4

/28

11111111.11111111.11111111.11110000

A 200.0.0.1 to 200.0.0.62

B 200.0.0.65 to 200.0.0.94

C 200.0.0.97 to 200.0.0.126

D 200.0.0.129 to 200.0.0.158

E 200.0.0.161 to 200.0.0.174


if you have any doubt then please ask me without any hesitation in the comment section below , if you like my answer then please thumbs up for the answer , before giving thumbs down please discuss the question it may possible that we may understand the question different way and we can edit and change the answers if you argue, thanks :)